A nurse is preparing to administer piperacillin 3.375 g by intermittent IV bolus every 6 hr. Available in piperacillin 3.375 g in dextrose 5% in water (D5W) 100 mL to infuse over 30 min. The nurse should set the IV pump to deliver how many mL/hr?
Answer #1

ANSWER.

* Calculation of ml/hr.

        ml/hr = Volume(ml) / Time(hr).

     Calculation of volume(ml).

          Volume(ml) = Desired dose/Dose in hand * Quantity.

   Here

         Desired dose (D) = 3.375 g.

         Dose in hand (H) = 3.375 g.

              Quantity (Q) = 100 ml.

    As per above formula

            Volume(ml) of piperacillin required = D/H * Q

                                                              = 3.375/3.375 * 100

                                                              = 100 ml.

                               Time(hr) = 0.5 hr (30 min)
          So

                 ml/hr = 100 / 0.5

                           = 200 ml/hr.